a. Give an estimated length for the P-P bond if the CI-cl bond is 200pm and a P-Cl bond is P-ci bond is 209pm.
Answer #1

8.

a. The half of the distance between the two nuclei of bonded atoms is called atomic radius. so the atomic radius of Cl-atom = 200 / 2

         = 100 pm

The atomic radius Cl-atom = 100 pm

By subtraction the atomic radius of Cl-atom from the P-Cl bond length, which gives atomic radius of P-atom,

The atomic radius P-atom = 209 - 100

                                                = 109 pm

The atomic radius of P-atom = 109 pm

The double of the atomic radius of P-atom is equal the P-P bond length,

P-P bond length = 2 * 109 pm

                              = 218 pm

The estimated bond length for the P-P bond = 218 pm